OpenStudy (anonymous):

Find the area of the region bounded by f(y) and g(y): f(y)=y^2 - 1 g(y)=4-(y-1)^2

OpenStudy (amistre64):

whats our points of intersection?

OpenStudy (anonymous):

Well, we need to set them equal to find them, right? Well, I don't really know how to get the exact numbers. I used the calculator and got: (.381966,-.618034) and (2.618034,1.618034)

OpenStudy (anonymous):

 Looks something like that

OpenStudy (amistre64):

y^2 - 1 = 4-(y-1)^2 y^2 - 1 = 4-(y^2-2y+1) y^2 - 1 = 4-y^2+2y-1 2y^2-2y = 4 y^2-y +1/4= 2+1/4 y = 1/2 +- 3/2 sound about right?

OpenStudy (amistre64):

y=2,-1

OpenStudy (amistre64):

so,\[\int_{-1}^{2}f(y)-g(y)\ dy\] and if you get a negative result ignore the sign.

OpenStudy (anonymous):

Alright, I got 9, is that the correct answer?

OpenStudy (amistre64):

dunno, i got know idea if you did the integrations right

OpenStudy (amistre64):

they are pretty basic tho so im sure they aint that hard to mess up :)

OpenStudy (anonymous):

yup lol thanks =)